The South Park DVD has been a slow seller during recent months. An analysis of monthly demand shows:

Q = 5,000 - 160P

where Q is DVD sales and P is price.

How many DVDs could be sold at a $25 price? (2points)

Calculate the point price elasticity of demand at a price of $25.
Answer #1

Q = 5,000 - 160P

(i) When P = 25, Q = 5,000 - (160 x 25) = 5,000 - 4,000 = 1,000

(ii) Elasticity of demand = (dQ/dP) x (P/Q) = - 160 x (25 / 1,000) = - 4